# IOPO: Empowering LLMs with Complex Instruction Following via Input-Output Preference Optimization

### v2

@@ Abstract
In the realm of large language models (LLMs), the ability of models to accurately follow instructions is paramount as more agents and applications leverage LLMs for construction, where the complexity of instructions are rapidly increasing. However, on the one hand, there is only a certain amount of complex instruction evaluation data; on the other hand, there are no dedicated algorithms to improve the ability to follow complex instructions. To this end, this paper introduces Trace, a benchmark for improving and evaluating the complex instruction-following ability, which consists of 120K training data and 1K evaluation data. Furthermore, we propose IOPO (Input-Output Preference Optimization) alignment method which takes both input and output preference pairs into consideration, where LLMs not only rapidly align with response preferences but also meticulously explore the instruction preferences. Extensive experiments on both in-domain and out-of-domain datasets confirm the effectiveness of IOPO, showing 8.15%, 2.18% improvements on in-domain data and 6.29%, 3.13% on out-of-domain data compared to SFT and DPO respectively. Our code and dataset are released at https://github.com/AlibabaResearch/DAMO-ConvAI/tree/main/IOPO.
@@
Xinghua Zhang, Haiyang Yu, Cheng Fu, Fei Huang, Yongbin Lithanks: Corresponding author. Tongyi Lab, Alibaba Group 🖂: {zhangxinghua.zxh, yifei.yhy, fucheng.fuc, f.huang, shuide.lyb}@alibaba-inc.com
@@ 1 Introduction
The rapid development of large language models (LLMs) has facilitated human-machine interaction, with instructions serving as the medium Gao et al. (2024); Kim et al. (2024); Zhang et al. (2024b). As human needs evolve, there is an increasing expectation for models to handle more intricate tasks through complex instructions Ge et al. (2023); Yang et al. (2024b); Wang et al. (2024). Consequently, the instruction-following ability, especially complex instructions, is garnering significant attention Zhou et al. (2023); Xu et al. (2024); Li et al. (2024); Zhang et al. (2024a).
@@
To evaluate the instruction-following abilities of LLMs, several benchmarks Zhou et al. (2023); Qin et al. (2024); Li et al. (2024) have been proposed, which are designed to systematically assess how well these models can understand and execute instructions. IFEval Zhou et al. (2023) focuses on verifiable instructions which are amenable to objective verification of compliance. Qin et al. (2024) introduces InFoBench which contains 500 distinct instructions and 2,250 decomposed questions to assess the ability of instruction following. Recently, the ability to follow complex instructions with multiple constraints is gaining increasing attention He et al. (2024b); Jiang et al. (2024); Wen et al. (2024); He et al. (2024a) as LLMs are deployed in sophisticated real-world applications. Zhang et al. (2024a) proposes a constraint-following benchmark CFBench with 1,000 multi-constraint samples. However, most of benchmarks lay emphasis on evaluating LLMs’ ability to follow complex instructions, lack of algorithms tailored for enhancing the corresponding ability.
@@
To bridge this gap, this paper first introduces Trace benchmark to improve the ability of LLMs to track complex fine-grained constraint instructions and make them more obedient. Trace is automatically constructed based on the manually sorted taxonomy of complex instructions with 26 constraint dimensions within 5 constraint types. We develop an automated data construction workflow that extends from open-source simple instructions to multi-constrained complex ones. In the end, we accumulate 120K complex instructions for model training and 1K human-verified data for evaluation. To enhance the ability of LLMs to follow complex instructions, this paper further proposes Input-Output Preference Optimization (IOPO) method. IOPO not only takes the instruction $x$ as input to directly learn the response $y$ preference, but also gradually delves deeper into instructions $x$ based on the same response $y$, to promote effective perception of fine-grained constraints, as shown in Figure 3 (b).
The major contributions of this paper are summarized as follows:
@@
  • Extensive experiments on both in-domain and out-of-domain evaluations have confirmed the consistent improvements, with an average increase of 7.22% and 2.66%, compared to SFT and DPO, respectively.
@@ 2.2 LLM Alignment
LLM alignment aims to enhance LLMs by aligning them with human preference. Recent research has conducted extensive explorations for LLM alignment. From RLHF/PPO Ouyang et al. (2022); Bai et al. (2022a) to DPO Rafailov et al. (2023) and beyond Bai et al. (2022b); Song et al. (2024); Meng et al. (2024), the evolution of xPO-series alignment algorithms has seen significant advancements. These methodologies have been pivotal in improving the alignment between LLMs and human values, ensuring that the outputs of these models are not only effective but also follow human preference.
RLHF involves training models using human-provided rewards or reward models to improve decision-making, optimizing policies through iterative feedback loops for more aligned outcomes. DPO directly optimizes the model’s output to match preferred response as indicated by human feedback, which simplifies the alignment process by focusing on direct comparisons between preferred and dispreferred outputs, allowing the model to learn human preference without needing an explicitly defined reward model. SimPO Meng et al. (2024) proposes a method for preference optimization that eliminates the need for the reference model $\pi_{\rm ref}$, which is memory efficient and simple. Instead of using pairwise data, PRO Song et al. (2024) utilizes the preference ranking of any length listwise preference dataset.
@@ 3 Preliminary
Figure 2: Construction Pipeline of Trace.
@@ 4.1 Construction Pipeline
The overall construction process includes several key stages: 1) Taxonomy of Constraint, 2) Constraint Expansion, 3) Instruction Structuring, 4) Quality Control, and 5) Response Generation & Evaluation.
@@ 4.3 Evaluation Protocol
We use GPT-4o as the evaluator to assess the generated response based on the complex instruction. Concretely, inspired by Qin et al. (2024); Zhang et al. (2024a), we prompt the LLM evaluator to evaluate each constraint mentioned in the complex instruction on a scale of 0 to 10 score, assessing the degree to which the response follows each constraint. A higher score indicates stronger adherence to the specified constraint. The overall instruction following score IF on the evaluation set with $n$ complex instructions can be calculated as follows:
@@ 4.4 Evaluation Set Quality
To generate the high-quality evaluation set, we further introduce a rigorous post-inspection process after construction pipeline (Sec. 4.1). First, we use the powerful LLM GPT-4o to check the following items for each instruction in the evaluation set: 1) Is the description empty? 2) Is there redundancy between the constraints and description? 3) Does the input match the description? If any of the aforementioned issues arise, we prompt the GPT-4o to make corrections. Second, we make the manual annotation process involving multiple steps such as annotator training, small-scale trial annotation, selection of official annotators, and formal annotation. Finally, we randomly select 100 instructions for quality evaluation, which are then inspected by three labeling specialists based on the above check items and the overall validity. The agreement rate among three annotators on the sampled evaluation set is 95%.
@@ 5 Input-Output Preference Optimization
Concretely, we construct a pair of instructions <$x_{1}$, $x_{2}$> whose responses are respectively <$y_{1}$, $y_{2}$>, where $x_{2}$ has subtle differences from $x_{1}$ in some constraints, and these differences would result in substantially divergent responses. And then, we can get four input-output pairs <$x_{1}$, $y_{1}$>, <$x_{1}$, $y_{2}$>, <$x_{2}$, $y_{1}$>, and <$x_{2}$, $y_{2}$>, which can form a preference group pair $\mathcal{G}_{1}$ $\succ$ $\mathcal{G}_{2}$ ($\mathcal{G}_{1}$ = {<$x_{1}$, $y_{1}$>, <$x_{2}$, $y_{2}$>}, $\mathcal{G}_{2}$ = {<$x_{1}$, $y_{2}$>, <$x_{2}$, $y_{1}$>}). The detailed data construction process is described in Appendix D. The first group is the matched input-output pair while the second one is mismatched. As derived from Eq.2 in Rafailov et al. (2023), the reward function $r(x,y)$ can be represented by the policy model $\pi_{r}$ as follows:
@@
Model Method Trace IFEval CFBench
IF-S IF-M S-Acc L-Acc CSR ISR PSR
Qwen2-7B Instruct 72.5 54.5 51.6 56.4 75.8 39.1 50.2
SFT 76.0 56.1 52.3 54.2 77.8 40.4 52.9
PPO 77.0 57.7 51.4 53.8 76.2 38.8 50.6
DPO 79.0 67.2 52.7 58.2 80.0 45.1 57.9
IOPO (Ours)Improv. 82.0↑3.0 68.9↑1.7 59.9↑7.2 63.6↑5.4 80.7↑0.7 47.0↑1.9 58.7↑0.8
Llama3.1-8B Instruct 67.5 52.9 74.3 78.6 71.4 35.7 46.9
SFT 75.5 62.9 71.0 74.1 78.4 43.2 54.7
PPO 75.0 57.3 69.9 72.3 75.9 40.9 50.7
DPO 79.0 69.2 71.5 76.5 80.8 48.1 59.8
IOPO (Ours)Improv. 81.5↑2.5 70.7↑1.5 78.2↑6.7 81.0↑4.5 81.8↑1.0 49.9↑1.8 61.1↑1.3
Table 3: Main results on in-domain Trace, and out-of-domain IFEval, and CFBench.
@@
Next, combining Eq. 6 and Eq. 4, we can further derive as follows (details in Appendix E):
@@
$\mathcal{L}$ (${}_{\rm IOPO}(\pi_{\theta})=-\mathbb{E}_{x_{1},y_{1},x_{2},y_{2}\sim D}\,\bigg{\{}{\rm log}\bigg{[}\sigma\bigg{(}\frac{1}{2}(2\beta{\rm log}\frac{\pi_{\theta}(y_{1}|x_{1})}{\pi_{\rm ref}(y_{1}|x_{1})}$ (9)
$-\beta{\rm log}\frac{\pi_{\theta}(y_{2}|x_{1})}{\pi_{\rm ref}(y_{2}|x_{1})}-\beta{\rm log}\frac{\pi_{\theta}(y_{1}|x_{2})}{\pi_{\rm ref}(y_{1}|x_{2})}+2\beta{\rm log}\frac{\pi_{\theta}(y_{2}|x_{2})}{\pi_{\rm ref}(y_{2}|x_{2})}$
$-\beta{\rm log}\frac{\pi_{\theta}(y_{1}|x_{2})}{\pi_{\rm ref}(y_{1}|x_{2})}-\beta{\rm log}\frac{\pi_{\theta}(y_{2}|x_{1})}{\pi_{\rm ref}(y_{2}|x_{1})})\bigg{)}\bigg{]}\bigg{\}}$
Model Method Trace IFEval CFBench
IF-S IF-M S-Acc L-Acc CSR ISR PSR
Qwen2-7B IOPO 82.0 68.9 59.9 63.6 80.7 47.0 58.7
w/o Output Pref 81.0 66.7 55.1 60.5 79.4 46.6 56.3
w/o Input Pref 80.9 67.1 56.7 61.9 79.7 46.8 57.0
Llama3.1-8B IOPO 81.5 70.7 78.2 81.0 81.8 49.9 61.1
w/o Output Pref 81.5 69.6 77.3 80.6 80.6 48.6 58.4
w/o Input Pref 79.0 69.0 77.9 80.2 80.9 48.3 59.4
Table 4: Ablation studies on Trace, IFEval, and CFBench.
@@ 6.1 Experimental Settings
Evaluation Datasets. We conduct experiments on three instruction-following datasets: Trace, IFEval Zhou et al. (2023), and CFBench Zhang et al. (2024a). Trace evaluation set is introduced in this paper, which has 1,042 instructions, and an average of 4.89 constraints per instruction, with a maximum of 15 constraints. IFEval consists of 541 prompts, with each prompt containing one or multiple verifiable instructions. CFBench contains 1,000 samples that cover more than 200 real-life scenarios and over 50 NLP tasks, with each sample including multiple constraints. It is worth noting that Trace is the in-domain evaluation set, IFEval and CFBench are the out-of-domain ones.
Implementation Details. (1) Trace Benchmark: we choose Qwen2-72B-Instruct Yang et al. (2024a) [^1] https://www.modelscope.cn/models/Qwen/Qwen2-72B-Instruct for benchmark construction. (2) IOPO Alignment: we choose Qwen2-7B-Instruct [^2] https://modelscope.cn/models/Qwen/Qwen2-7B-Instruct, and LLaMA3.1-8B-Instruct [^3] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct as the LLM backbone. All models, except for the base models (i.e. Qwen2-7B-Instruct and Llama-3.1-8B-Instruct), are trained on Trace’s training set or its variants (Train Once, Test Anywhere). The learning rate is 1e-4 for supervised fine-tuning (SFT), and 5e-6 for DPO and IOPO. The maximum length and epoch are set to 6,000 and 3 respectively. $\beta$ is set to 0.1. We implement our code based on LLaMA-Factory Zheng et al. (2024), perform parallel training on 4 $\times$ 8-GPU machines, with a batch size of 1 per GPU. The DPO training data construction is shown in Appendix C.
Evaluation Metrics. For Trace, we use GPT-4o to evaluate if all constraints in the instruction have been followed (IF-S for single-constraint instructions, and IF-M for multi-constraint instructions), as described in Sec. 4.3. For IFEval, we use prompt-level strict and loose accuracy defined in Zhou et al. (2023), abbr. S-Acc and L-Acc respectively. CFBench Zhang et al. (2024a) introduces three evaluation metrics with GPT-4o as the evaluation model: constraint satisfaction rate (CSR), instruction satisfaction rate (ISR), and priority satisfaction rate (PSR).

### v3

@@ Abstract
In the realm of large language models (LLMs), the ability of models to accurately follow instructions is paramount as more agents and applications leverage LLMs for construction, where the complexity of instructions are rapidly increasing. However, on the one hand, there is only a certain amount of complex instruction evaluation data; on the other hand, there are no dedicated algorithms to improve the ability to follow complex instructions. To this end, this paper introduces Trace, a benchmark for improving and evaluating the complex instruction-following ability, which consists of 120K training data and 1K evaluation data. Furthermore, we propose IOPO (Input-Output Preference Optimization) alignment method which takes both input and output preference pairs into consideration, where LLMs not only rapidly align with response preferences but also meticulously explore the instruction preferences. Extensive experiments on both in-domain and out-of-domain datasets confirm the effectiveness of IOPO, showing 8.15%, 2.18% improvements on in-domain data and 5.91%, 2.83% on out-of-domain data compared to SFT and DPO respectively. Our code and dataset are released at https://github.com/AlibabaResearch/DAMO-ConvAI/tree/main/IOPO.
@@
Xinghua Zhang, Haiyang Yu, Cheng Fu, Fei Huang, Yongbin Lithanks: Corresponding author. Tongyi Lab, Alibaba Group {zhangxinghua.zxh, yifei.yhy, fucheng.fuc, f.huang, shuide.lyb}@alibaba-inc.com
@@ 1 Introduction
The rapid development of LLMs has facilitated human-machine interaction, with instructions serving as the medium Gao et al. (2024); Kim et al. (2024); Zhang et al. (2024c); Wang et al. (2024b); Zhang et al. (2024b). As human needs evolve, there is an increasing expectation for models to handle more intricate tasks through complex instructions Ge et al. (2023); Yang et al. (2024b); Wang et al. (2024a). Consequently, the instruction-following ability, especially complex instructions, is garnering significant attention Zhou et al. (2023); Xu et al. (2024); Li et al. (2024); Zhang et al. (2024a).
@@
To evaluate the instruction-following abilities of LLMs, several benchmarks Zhou et al. (2023); Qin et al. (2024); Li et al. (2024) have been proposed, which are designed to systematically assess how well these models can understand and execute instructions. IFEval Zhou et al. (2023) focuses on verifiable instructions which are amenable to objective verification of compliance. Qin et al. (2024) introduces InFoBench which contains 2,250 decomposed questions to assess the instruction following. Recently, the ability to follow complex instructions with multiple constraints is gaining increasing attention He et al. (2024b); Jiang et al. (2024); Wen et al. (2024); He et al. (2024a) as LLMs are deployed in sophisticated real-world applications. Zhang et al. (2024a) proposes a constraint-following benchmark CFBench with 1,000 multi-constraint samples. However, most of benchmarks lay emphasis on evaluating LLMs’ ability to follow complex instructions, lack of algorithms tailored for enhancing the corresponding ability.
@@
To bridge this gap, this paper first introduces Trace benchmark to improve the ability of LLMs to track complex fine-grained constraint instructions and make them more obedient. Trace is automatically constructed based on the manually sorted taxonomy of complex instructions with 26 constraint dimensions within 5 constraint types. We develop an automated data construction workflow that extends from open-source simple instructions to multi-constrained complex ones. In the end, we accumulate 120K complex instructions for model training and 1K human-verified data for evaluation. To enhance the ability of LLMs to follow complex instructions, this paper further proposes Input-Output Preference Optimization (IOPO) method. IOPO not only takes the instruction $x$ as input to directly learn the response $y$ preference, but also gradually delves deeper into instructions $x$ based on the same response $y$, to promote effective perception of fine-grained constraints, as shown in Figure 1 (b). The major contributions of this paper are summarized as follows:
@@
  • Extensive experiments on both in-domain and out-of-domain evaluations have confirmed the consistent improvements, with an average increase of 7.03% and 2.51%, compared to SFT and DPO, respectively.
@@ 2.2 LLM Alignment
LLM alignment aims to enhance LLMs by aligning them with human preference. Recent research has conducted extensive explorations for LLM alignment. From RLHF/PPO Ouyang et al. (2022); Bai et al. (2022a) to DPO Rafailov et al. (2023) and beyond Bai et al. (2022b); Song et al. (2024); Meng et al. (2024), the evolution of xPO-series alignment algorithms has seen significant advancements. These methods have been pivotal in improving the alignment between LLMs and human values, ensuring that the outputs of these models are not only effective but also follow human preference.
RLHF involves training models using human-provided rewards or reward models to improve decision-making, optimizing policies through iterative feedback loops for more aligned outcomes. DPO directly optimizes the model’s output to match preferred response as indicated by human feedback, which simplifies the alignment process by focusing on direct comparisons between preferred and dispreferred outputs, allowing the model to learn human preference without needing an explicitly defined reward model. SimPO Meng et al. (2024) proposes a method for preference optimization that eliminates the need for the reference model $\pi_{\rm ref}$, which is memory efficient and simple. Instead of using pairwise data, PRO Song et al. (2024) utilizes the preference ranking of any length listwise preference dataset. ORPO Hong et al. (2024) introduces the reference model-free preference optimization algorithm.
@@ 3 Preliminary
Figure 2: Construction Pipeline of Trace, which consists of five key stages: 1) Taxonomy of Constraint, 2) Constraint Expansion, 3) Instruction Structuring, 4) Quality Control, and 5) Response Generation & Evaluation.
@@ 4.1 Construction Pipeline
The overall construction process as shown in Figure 2 includes several key stages: 1) Taxonomy of Constraint, 2) Constraint Expansion, 3) Instruction Structuring, 4) Quality Control, and 5) Response Generation & Evaluation.
@@ 4.3 Evaluation Protocol
Following previous work Qin et al. (2024); Zhang et al. (2024a), we use GPT-4o as the evaluator to assess the generated response based on the complex instruction. Concretely, we prompt the LLM evaluator to evaluate each constraint mentioned in the complex instruction on a scale of 0–10, assessing the degree to which the response follows each constraint. A higher score indicates stronger adherence to the specified constraint. The overall instruction following score IF on the evaluation set with $n$ complex instructions are calculated as follows:
@@ 4.4 Evaluation Set Quality
To generate the high-quality evaluation set, we further introduce a rigorous post-inspection process after the construction pipeline (Sec. 4.1). First, based on LLM-as-Judge Wang et al. (2024c); Zhang et al. (2023); Zeng et al. (2024); Xia et al. (2024), we use the powerful LLM GPT-4o to check the following items for each instruction in the evaluation set: 1) Is the description empty? 2) Is there redundancy between the constraints and description? 3) Does the input match the description? If any of the aforementioned issues arise, we prompt the GPT-4o to make corrections for accelerating the subsequent manual verification. Second, the professional data annotation team makes the manual annotation process involving multiple steps such as annotator training, small-scale trial annotation, selection of official annotators, and formal annotation. Finally, we randomly select 100 instructions for quality evaluation, which are then inspected by three labeling specialists based on the above check items and the overall validity. The agreement rate among three annotators on the sampled evaluation set is 95%.
@@ 5 Input-Output Preference Optimization
Concretely, we construct a pair of instructions <$x_{1}$, $x_{2}$> whose responses are respectively <$y_{1}$, $y_{2}$>, where $x_{2}$ has subtle differences from $x_{1}$ in some constraints, and these differences would result in substantially divergent responses. And then, we can get four input-output pairs <$x_{1}$, $y_{1}$>, <$x_{1}$, $y_{2}$>, <$x_{2}$, $y_{1}$>, and <$x_{2}$, $y_{2}$>, which can form a preference group pair $\mathcal{G}_{1}$ $\succ$ $\mathcal{G}_{2}$ ($\mathcal{G}_{1}$ = {<$x_{1}$, $y_{1}$>, <$x_{2}$, $y_{2}$>}, $\mathcal{G}_{2}$ = {<$x_{1}$, $y_{2}$>, <$x_{2}$, $y_{1}$>}). The detailed data construction process is described in Appendix D. The first group is the matched input-output pair while the second one is mismatched. As derived in Rafailov et al. (2023), the reward function $r(x,y)$ can be represented by the policy model $\pi_{r}$ as follows:
@@
Next, combining Eq. 6 and Eq. 4, we can further derive as follows (details in Appendix G):
@@
$\mathcal{L}_{\rm IOPO}(\pi_{\theta})$ $=-\mathbb{E}_{i\sim D}\,\bigg{\{}{\rm log}\bigg{[}\sigma\bigg{(}\frac{\Pi_{1}(\pi_{\theta})+\Pi_{2}(\pi_{\theta})}{2}\bigg{)}\bigg{]}\bigg{\}}$ (9)
$i$ $=<x_{1},y_{1},x_{2},y_{2}>$
$\Pi_{1}(\pi_{\theta})$ $=\bigg{(}\underbrace{\beta\log\frac{\pi_{\theta}(y_{1}|x_{1})}{\pi_{\rm ref}(y_{1}|x_{1})}-\beta\log\frac{\pi_{\theta}(y_{2}|x_{1})}{\pi_{\rm ref}(y_{2}|x_{1})}}_{\rm\mathtt{Output}}\bigg{)}$ (10)
$+\bigg{(}\underbrace{\beta\log\frac{\pi_{\theta}(y_{1}|x_{1})}{\pi_{\rm ref}(y_{1}|x_{1})}-\beta\log\frac{\pi_{\theta}(y_{1}|x_{2})}{\pi_{\rm ref}(y_{1}|x_{2})}}_{\rm\mathtt{Input}}\bigg{)}$
$\Pi_{2}(\pi_{\theta})$ $=\bigg{(}\underbrace{\beta\log\frac{\pi_{\theta}(y_{2}|x_{2})}{\pi_{\rm ref}(y_{2}|x_{2})}-\beta\log\frac{\pi_{\theta}(y_{1}|x_{2})}{\pi_{\rm ref}(y_{1}|x_{2})}}_{\rm\mathtt{Output}}\bigg{)}$
$+\bigg{(}\underbrace{\beta\log\frac{\pi_{\theta}(y_{2}|x_{2})}{\pi_{\rm ref}(y_{2}|x_{2})}-\beta\log\frac{\pi_{\theta}(y_{2}|x_{1})}{\pi_{\rm ref}(y_{2}|x_{1})}}_{\rm\mathtt{Input}}\bigg{)}$
where we mark the preference modeling for $\mathtt{Output}$ and $\mathtt{Input}$ in $\Pi_{1}(\pi_{\theta})$, $\Pi_{2}(\pi_{\theta})$.
Model Method Trace IFEval CFBench ComplexBench
$\mathtt{IF}$-$\mathtt{S}$ $\mathtt{IF}$-$\mathtt{M}$ $\mathtt{S}$-$\mathtt{Acc}$ $\mathtt{L}$-$\mathtt{Acc}$ $\mathtt{CSR}$ $\mathtt{ISR}$ $\mathtt{PSR}$ $\mathtt{DRFR}$
Qwen2-7B Instruct 72.5 54.5 51.6 56.4 75.8 39.1 50.2 68.1
SFT 76.0 56.1 52.3 54.2 77.8 40.4 52.9 68.2
PPO 77.0 57.7 51.4 53.8 76.2 38.8 50.6 68.6
ORPO 77.9 61.7 53.1 56.9 79.7 45.9 57.0 69.1
SimPO 78.3 63.6 52.2 57.6 78.4 45.0 57.6 67.8
DPO 79.0 67.2 52.7 58.2 80.0 45.1 57.9 70.9
IOPO (Ours)Improv. 82.0↑3.0 68.9↑1.7 59.9↑7.2 63.6↑5.4 80.7↑0.7 47.0↑1.9 58.7↑0.8 72.6↑1.7
Llama3.1-8B Instruct 67.5 52.9 74.3 78.6 71.4 35.7 46.9 63.2
SFT 75.5 62.9 71.0 74.1 78.4 43.2 54.7 68.2
PPO 75.0 57.3 69.9 72.3 75.9 40.9 50.7 68.1
ORPO 77.0 63.1 72.3 77.3 79.4 46.6 57.2 69.6
SimPO 76.3 64.5 71.2 76.6 80.6 47.8 58.7 70.2
DPO 79.0 69.2 71.5 76.5 80.8 48.1 59.8 70.8
IOPO (Ours)Improv. 81.5↑2.5 70.7↑1.5 78.2↑6.7 81.0↑4.5 81.8↑1.0 49.9↑1.8 61.1↑1.3 71.8↑1.0
Table 3: Main results on in-domain Trace, and out-of-domain IFEval, CFBench, and ComplexBench. ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}Improv.}$ indicates the absolute improvement compared to DPO.
Model Method Trace IFEval CFBench ComplexBench
$\mathtt{IF}$-$\mathtt{S}$ $\mathtt{IF}$-$\mathtt{M}$ $\mathtt{S}$-$\mathtt{Acc}$ $\mathtt{L}$-$\mathtt{Acc}$ $\mathtt{CSR}$ $\mathtt{ISR}$ $\mathtt{PSR}$ $\mathtt{DRFR}$
Qwen2-7B IOPO 82.0 68.9 59.9 63.6 80.7 47.0 58.7 72.6
w/o $\mathtt{Output}$ $\mathtt{Preference}$ 81.0 66.7 55.1 60.5 79.4 46.6 56.3 71.0
w/o $\mathtt{Input}$ $\mathtt{Preference}$ 80.9 67.1 56.7 61.9 79.7 46.8 57.0 71.3
Llama3.1-8B IOPO 81.5 70.7 78.2 81.0 81.8 49.9 61.1 71.8
w/o $\mathtt{Output}$ $\mathtt{Preference}$ 81.5 69.6 77.3 80.6 80.6 48.6 58.4 69.2
w/o $\mathtt{Input}$ $\mathtt{Preference}$ 79.0 69.0 77.9 80.2 80.9 48.3 59.4 70.1
Table 4: Ablation studies on Trace, IFEval, CFBench, and ComplexBench.
@@ 6.1 Experimental Settings
Evaluation Datasets. We conduct experiments on four instruction-following datasets: Trace, IFEval Zhou et al. (2023), CFBench Zhang et al. (2024a), and ComplexBench Wen et al. (2024). Trace evaluation set is introduced in this paper, which has 1,042 instructions, and an average of 4.89 constraints per instruction, with a maximum of 15 constraints. IFEval consists of 541 prompts, with each prompt containing one or multiple verifiable instructions. CFBench contains 1,000 samples that cover more than 200 real-life scenarios and over 50 NLP tasks, with each sample including multiple constraints. ComplexBench creates 1,150 samples based on 4 constraint types, 19 constraint dimensions, and 4 composition types. It is worth noting that Trace is the in-domain evaluation set, IFEval, CFBench and ComplexBench are the out-of-domain ones.
Implementation Details. (1) Trace Benchmark: we choose Qwen2-72B-Instruct Yang et al. (2024a)111 https://modelscope.cn/models/Qwen/Qwen2-72B-Instruct for benchmark construction. (2) IOPO Alignment: we choose Qwen2-7B-Instruct [^2] https://modelscope.cn/models/Qwen/Qwen2-7B-Instruct, and LLaMA3.1-8B-Instruct [^3] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct as the LLM backbone. All models, except for the base models (i.e. Qwen2-7B-Instruct and Llama-3.1-8B-Instruct), are trained on Trace’s training set or its variants, tested on all evaluation datasets (Train Once, Test Anywhere). The learning rate is 1e-4 for supervised fine-tuning (SFT), and 5e-6 for DPO and IOPO. The maximum length and epoch are set to 6,000 and 3 respectively. $\beta$ is set to 0.1. We implement our code based on LLaMA-Factory Zheng et al. (2024), perform parallel training on 4 $\times$ 8-GPU machines, with a micro batch size of 1 per GPU. The DPO training data construction is shown in Appendix C.
Evaluation Metrics. For Trace, we use GPT-4o to evaluate if all constraints in the instruction have been followed (IF-S for single-constraint instructions, and IF-M for multi-constraint instructions), as described in Sec. 4.3. For IFEval, we use prompt-level strict and loose accuracy defined in Zhou et al. (2023), abbr. S-Acc and L-Acc respectively. CFBench Zhang et al. (2024a) introduces three evaluation metrics with GPT-4o as the evaluation model: constraint satisfaction rate (CSR), instruction satisfaction rate (ISR), and priority satisfaction rate (PSR). ComplexBench Wen et al. (2024) calculates the decomposed requirements following ratio (DRFR).